In a group of 20 batteries, 3 are dead. You choose 2 batteries at random a) Create a probability model for the number of good batteries you get. b) Whats the expected number of good ones you get? c) Whats the standard deviation? a) Create a probability model. Number good 0 2 P(Number good) (Round to three decimal places as needed.)can someone show me how to do this thank you please i inned inneeeed it
Answer #1

a)

Let X = Number of batteries are good.

For X = 0 , That is both batteries are defective.

Number of ways to get 2 defective batteries from 3 defective = 3C2

Total number of ways to get 2 batteries from 20 =  20C2

P( X = 0) =  3C2 / 20C2

= 3 / 190

= 0.016

For X = 1

That is we choose 1 out of 2 batteries is good.

P(1 good battery) = P( one good) * P(one defective)

= 17C1 * 3C1 / 20C2

= 0.268

For X = 2

That is we choose both good batteries from 17 good batteries.

P( Both good) = 17C2 / 20C2

= 136 / 190

= 0.716

Number of good 0    1 2
Probability of good 0.016 0.268 0.716

b)

Let X be number of good batteries

P(X) be probability of good batteries

E(X) = \sum X * P(X)

= 0 * 0.016 + 1* 0.268 + 2 * 0.716

= 1.7

Expected number of good batteries = 1.7

c)

Standard deviation = Sqrt [\sum X2 * P(X) - [E(X) ]2 ]

= Sqrt[ (02 * 0.016 + 12 * 0.268 + 22 * 0.716 ) - 1.72 ]

= 0.4919